PUBLICATION LIST

1. Cheng, G., Mills, K., Podgorny, M. and Fox, G.C., “Developing Interactive Parallel Programs on IBM SP-1 within AVS Framework,”, submitted to AVS ’94 Conference, Boston, MA, May 2-4, 1994.

2. Podgorny, M., Cheng, G., Markowski, R., Furmanski, W., Hariri, S. and Fox, G.C., “Using AVS to Implement Heterogenous Distributed Applications on NPAC Metacomputer,” submitted to AVS ’94 Conference, Boston, MA, May 2-4, 1994.

3. Cheng, G., Fox, G.C. and Hawick, K., “A Scaleable Paradigm for Effectively-Dense Matrix Formulated Applications,” submitted to the European Conference and Exhibition on High-Performance Computing and Networking (HPCN Europe) 1994, Munich, Germany, April 18-20, 1994.

4. Cheng, G., Mills, K. and Fox, G.C., ”Integrating PVM-based Network Concurrent Computing into Application Visualization System(AVS) Framework,” submitted to the 1994 Scalable High Performance Computing Conference, Knoxville, Tennessee, May 23-25, 1994.

5. Seablom, M., Lyster P., Mills, K., Laszewski, G. and Cheng, G., “The NASA Data Analysis System for Massively Parallel Computers,” submitted to 1994 SCS Simulation Multiconference, High Performance Computing ’94, April 11-15, 1994, La Jolla, CA.

6. Mills, K., Shen, X., Cheng, G., Miller, D., “A Massive Data Parallel Program for 3D Proton Dose Distribution,” submitted to 1994 SCS Simulation Multiconference, High Performance Computing ’94, April 11-15, 1994, La Jolla, CA.

7. Cheng, G., Lu, Y., Fox, G.C., Mills K. and Haupt, T., “An Interactive Remote Visualization Environment for an Electromagnetic Scattering Simulation on a High Performance Computing System,” to be published in Proc. of IEEE Supercomputing ‘93, Portland, OR, November 1993, IEEE Computer Society Press, Los Alamitos, CA.

8. Cheng, G., Sparling, L., Mills, K. and Lyster P., “Four Dimensional Data Assimilation (NASA HPCC): Parallel Trajectory Model,” Poster presentation at IEEE Supercomputing ‘93, Portland, OR, November 1993, IEEE Computer Society Press, Los Alamitos, CA.

9. Cheng, G., Faigle, C., Fox, G.C., Fumanski, W., Li, B. and Mills, K., “Exploring AVS for HPDC Software Integration: Case Studies Towards Parallel Support for GIS,” Proc. of the 2nd AVS Conference AVS’93, Lake Buena Vista, FL, May 1993.

10. Cheng, G., Mills, K., Fox, G.C., “An Interactive Visualization Environment for Financial Modeling on Heterogeneous Computing Systems,” in Proc. of the 6th SIAM Conference on Parallel Processing for Scientific Computing, March 1993, Norfolk, VA, pp. 891-989, Vol. 2.

11. Li, B., Mills, K., Csillag, F., Podgorny, M., Vaughn, B. and Cheng, G., “Integration of GRASS with High Performance Parallel Computing Environment, “ in Proc. of the 8th GRASS GIS Users Conference, Open Grass Foundation, Reston, VA, March 1993.

12. Mills, K., Cheng, G., Vinson, M., Ranka, S., and Fox, G.C., “Software Issues and Performance of a Stock Option Pricing Model on the Connection Machine-2 and DECmpp-12000,” in Proc. of Fifth Australian Supercomputing Conference, World Congress Centre, Melbourne, Australia, December, 1992, pp. 125-134.

13. Mills, K., Vinson, M., Cheng, G., “A Large Scale Comparison of Option Pricing Models with Historical Market Data,” in Proceedings of 4th Symposium on the Frontiers of Massively Parallel Computation (McLean, Virginia, October 19-21, 1992), IEEE Computer Society Press, Los Alamitos, California, 1992, pp. 420-426.

14. Cheng, G. and Zhang, Y., “A Functional + Logic Programming Language in Interpretation-Compilation Implementation,” Lisp And Symbolic Computation: An International Journal, Vol. 5, No. 3, 1992, pp. 133-156.

15. Cheng, G., Huang, S., et al, “An Employee Attendance System with Mag-Card Reader,” Microcomputer Applications, pp. 15-18, Vol. 12, No. 4, July, 1991(in Chinese).

16. Huang, X. and Cheng, G., “A Province-Wide Natural Disaster Precaution System: An Application of Geographic Information System to Natural Disaster Reduction,” in Proc. of the 2nd International Workshop on GIS, Beijing, pp. 386-391, 1990.

17. Cheng, G. and Zhang, Y., “A Functional Plus Logic Programming Language,” Mini- and Micro- Computers, No. 1, Vol. 11, 1990(in Chinese), pp. 14-19.

18. Cheng, G., Huang, S., et al, “A Highway Toll-Collecting System with Mag-Card Reader,” Computer Applications, pp. 10-13, No. 6, 1990(in Chinese).

19. Cheng, G. and Kang, Z., “A Computer-Based Information System on Moving Elements of Debris Flow,” in Proc. of the 1st Japan-China Symposium on Landslides and Debris Flows, pp. 195-198, Tokyo, Japan, 1989.

20. Cheng, G. and Zhang, Y., “Designing A Knowledge-based Programming Environment with Functional and Logic Paradigms,” Computer Applications, pp. 20-25, No. 4, 1989(in Chinese).

21. Cheng, G., “The Implementation of An AI Language Unifying Functional and Logic Paradigms, M.S. Thesis,” Computer Science and Engineering Department, Huazhong University of Science and Technology, China, 1988, 77 pps (in Chinese).

22. Kang, Z., Hu, P., Cheng, G., “Sorting and Stockpile on Observed Data of Moving Elements of Debris Flow,” Mountain Research, pp. 289-294, Vol. 5, No. 4, Dec., 1987 (in Chinese).

23. Cheng, G., “Program Understanding,” Journal of HUST Graduates, pp. 52-55, No. 1, 1987 (in Chinese).

24. Cheng, G., “A New Programming Language TAO integrating LISP, Prolog and Smalltalk,” Journal of HUST Graduates, pp. 182-187, No. 1, 1988, and in Proc. of the 8th National Symposium on Artificial Intelligence, 1987 (in Chinese).

25. Mills, K., Cheng, G., Vinson, M., Ranka, S., and Fox, G.C., “Load Balancing, Communication, and Performance of a Stock Option Pricing Model on the Connection Machine-2 and DECmpp-12000,” Technical Report, Syracuse Center for Computational Science-273, 22 pps, March, 1992.

26. Cheng, G. and Fox, G. C., “An Interactive Graphic Tool for the Option Price Modeling on DECmpp-12000 Using AVS,” Technical Report, Syracuse Center for Computational Science-369,12 pps, August, 1992.

27. Cheng, G. and Fox, G. C., “CM-Fortran Implementation for American (Full) Binomial Approximation of a Stock Option Pricing Model,” Technical Report, Syracuse Center for Computational Science-225, 10 pps, December, 1991.

28. Cheng, G., Fox, G. C. and Mills K., “Integrating Multiple Programming Paradigms on Connection Machine CM5 in a Dataflow-based Software Environment,” Technical Report, Syracuse Center for Computational Science, 15 pps, August, 1993.